sarahms:

12P10/12P2 please help

BenLindquist:

what are you asking of?

sarahms:

Probability

BenLindquist:

P(12,10)/P(12,2) P(n,k)=n!/(n-k) P(12,10)/P(12,2)=(12!/2!)/(12!/10!) =10/2 =3628800/2 = 1,814,400 is this anywhere near it?

BenLindquist:

you didn't give us much of a choice to work through

sarahms:

all it says is to simplify the expression

BenLindquist:

look at what I wrote and see if that works.

sarahms:

thats it thank you!!!!

BenLindquist:

yw if you can please give me the best response